# General Lagos 1163, Valdivia 5110683, Chile

### v1

@@ 1 Introduction
Recently, it has been realized that it is possible to deform Maxwell theory with a dimensionless coupling $\gamma$, maintaining, at the same time, the electromagnetic duality of the dynamics and the invariance of the action under the conformal group Bandos:2020jsw . The latter implies that if such a continuous deformation exists, it must be controlled by a dimensionless coupling. Such a theory, discovered in Ref. Bandos:2020jsw , was dubbed ModMax. It was later rederived in Ref. Kosyakov:2020wxv using the Gaillard-Zumino-Gibbons-Rasheed formalism for duality rotations in nonlinear electrodynamics Gibbons:1995cv ; Gaillard:1981rj ; Gibbons:1995ap ; Gaillard:1997zr ; Gaillard:1997rt , and more recently in Ref. Ayon-Beato:2024vph employing the Plebański formulation.
@@
Since ModMax emerges naturally as the most general nonlinear electrodynamics possessing the same features as Maxwell theory, its physical consequences have been thoroughly explored over the last few years. The study of the spectrum of propagating degrees of freedom was done in Ref. Bandos:2020jsw , where it was also shown that there is birrefringence in vacuum. Supersymmetric extensions were constructed in Bandos:2021rqy , while in Flores-Alfonso:2020euz ; Amirabi:2020mzv ; Kubiznak:2022vft ; Barrientos:2022bzm ; Barrientos:2024umq it was shown that its coupling to gravity leads to interesting curved spacetimes as accelerating black holes, multi-black hole solutions Bokulic:2025usc , Euclidean solitons Corral:2025npd , NUTs and Bolts BallonBordo:2020jtw ; Flores-Alfonso:2020nnd , Eguchi-Hanson instantons Colipi-Marchant:2023awk , and hairy black holes Zhang:2021qga ; Ayon-Beato:2024vph . The Galilean and Carrollian limits have been taken and explored in Banerjee:2022sza ; Correa:2024qej (see also Chen:2024vho ). Recently, holographic aspects of ModMax theory have been explored in Ref. Barrientos:2025rde and their connection to $T\bar{T}$-deformations in conformal field theories was studied in Ref. Ferko:2022cix ; Babaei-Aghbolagh:2022leo ; Ferko:2023wyi .
@@
In Yang-Mills theory, instantons are Euclidean, stationary, and regular (anti-)self-dual solutions that minimize the action on a given topological sector Belavin:1975fg . They also control quantum effects as tunneling between topologically inequivalent vacua, and they are fundamental in a non-perturbative formulation of the quantum theory tHooft:1976rip ; tHooft:1976snw . However, due to to topological reasons, Euclidean $U(1)$ gauge theories do not have finite action, localized, regular instantonic solutions on flat space, a feature that is cured by the non-Abelian extensions. Given the relevance of the instantonic configurations, it is natural to ask whether they exist in a non-Abelian extension of the ModMax theory, where the notion of self-duality is natural, and how they are deformed with respect to the instantons in Yang-Mills. This is the aim of the present work.
@@
We start by considering a sensible one-parameter family of non-Abelian extension of ModMax theory in Sec. 2. The theory is parameterized by a single dimensionless constant $\gamma$, which reduces to Yang-Mills theory in the $\gamma\to 0$ limit. We demonstrate that the theory admits (anti-)self-dual configurations with a vanishing energy-momentum tensor, despite the inherent nonlinearities of the ModMax model. Focusing on the $SU(2)$ case, in Sec. 3 we construct the generalized BPST instanton on flat space and extend the solution to Euclidean de Sitter ($\mathbb{S}^{4})$ and anti-de Sitter ($\mathbb{H}^{4})$ backgrounds. On the latter case, the Chern-Pontryagin index is shown to depend on the instanton size, since the configuration is not a pure gauge at infinity; a feature already reported by Callan and Wilczek in Yang-Mills theory on negative-curvature backgrounds Callan:1989em . Section 4 considers the ’t Hooft ansatz, which accommodates multi-instantons in Yang-Mills theory tHooft:1976snw . Indeed, we show that such a choice is also very useful in the non-Abelian ModMax context. Remarkably, the resulting (anti-)self-dual system reduces to a single nonlinear equation, which we solve perturbatively in the dimensionless parameter. To first order, we obtain a formal solution for the $N$-instanton configuration. In Sec. 5, we compute the spectrum of the Dirac operator on the boundary of Euclidean AdS space, which is crucial for obtaining the non-local contributions to the Dirac index in the presence of non-Abelian fields. In Sec. 6, we couple the Euclidean non-Abelian ModMax theory to Einstein gravity in the presence of a conformally-coupled scalar field and construct new gravitating solutions, including Euclidean wormholes and gravitational instantons with a secondary scalar hair. We conclude with comments in Sec. 7.
@@ 2 Non-Abelian ModMax theory
The generalization of ModMax electrodynamics to the $SU(2)$ case has been studied in Ref. Cirilo-Lombardo:2023poc . It was shown that the latter can be obtained from the secular equation that arises from the four eigenvectors associated with the four main directions in the Newman-Penrose classification Newman:1961qr . The resulting theory is invariant under conformal and $SO(2)$ duality transformations and, in Euclidean signature, its action principle is given by [^1] Here, we will work mainly in Euclidean signature unless otherwise stated. The Lorentzian version of Eq. (4) is obtained by simultaneously changing the global sign of the action and performing $Y^{2}\to-Y^{2}$.
@@
It is worth mentioning that Gibbons and Rasheed studied $SO(2)$ electromagnetic duality invariance of nonlinear electrodynamics in Ref. Gibbons:1995cv . They concluded that the condition $\tilde{F}_{\mu\nu}F^{\mu\nu}=\tilde{P}_{\mu\nu}P^{\mu\nu}$ ensures that the theory is invariant under duality rotations. One can check that the non-Abelian generalization of the Gibbons-Rasheed condition, that is,
@@
is satisfied if the definition in Eq. (6) is assumed. Additionally, conformal invariance implies that the stress-energy tensor must be traceless. In the case of the non-Abelian theory, the latter is given by
@@
$F=\frac{1}{2}F_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}=\left(a^{\prime}\text{d}r\wedge\sigma^{i}+\frac{a(a-1)}{2}\epsilon^{i}_{\penalty 10000\ jk}\sigma^{b}\wedge\sigma^{c}\right)t_{i}\,$, (16)
@@
$P=\frac{1}{2}P_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}$ $=\left(e^{-\gamma}a^{\prime}\text{d}r\wedge\sigma^{i}+\frac{a(a-1)e^{\gamma}}{2}\epsilon^{i}_{\penalty 10000\ jk}\sigma^{j}\wedge\sigma^{k}\right)t_{i}\,$,
@@ 4 Beyond the BPST instanton
where $\Phi=\Phi(x)$ is a function of the coordinates, with no particular symmetry assumed, and $\bar{\eta}_{\mu\nu}:=\bar{\eta}_{a\mu\nu}\,t^{a}$. In the case of Yang-Mills theory, substituting this ansatz into the (anti-)self-dual condition $F_{\mu\nu}=\pm\tilde{F}_{\mu\nu}$ reduces them to $\nabla^{2}\Phi(x)=0$, which is solved by tHooft:1976snw
@@ 5 Dirac spectrum on the Euclidean AdS boundary
$\lambda_{\pm}$ $=\frac{1}{4}\pm 2\sqrt{\left(p-q-2a_{0\pm}\right)^{2}+4\left(2a_{0\pm}+pq\right)-8a_{0\pm}\sqrt{pq}}\;,$ (42)
$\lambda_{0}$ $=\frac{1}{4}+2\sqrt{\left(p-2a_{0\pm}\right)^{2}+8a_{0\pm}}\;,$ (43)
for $q\neq 0$ and $q=0$, respectively. Then, the contribution of non-Abelian ModMax instantons to the spectrum of the Dirac operator becomes transparent. This provides an important step towards obtaining the spectral asymmetry of the Dirac operator at the conformal boundary of Euclidean AdS, which is crucial for computing the index of the Dirac operator analytically. If the latter is nonvanishing, it would indicate the presence of axial anomaly induced by self-dual instantons in non-Abelian ModMax theory.
@@ 6.3 Euclidean AdS wormholes and gravitational instantons
Despite their Euclidean nature, they exhibit a wide range of applications in both particle physics and cosmology Hebecker:2018ofv . They also give rise to several conceptual challenges, such as the factorization problem in holography Maldacena:2004rf ; Arkani-Hamed:2007cpn ; Bergshoeff:2004pg ; Andriolo:2022rxc ; Jonas:2023ipa ; Gutperle:2002km . A broad variety of such configurations have been explored in different settings, including supergravity theories Hertog:2017owm ; Anabalon:2018rzq ; Anabalon:2020loe ; Anabalon:2023kcp ; Astesiano:2023iql , higher-curvature extensions of gravity Dotti:2006cp ; Dotti:2007az ; Dotti:2010bw , and in the presence of matter fields such as Yang-Mills fields Maldacena:2004rf ; Hosoya:1989zn , phantom fields Blazquez-Salcedo:2020nsa ; Chew:2021vxh , Higgs fields Betzios:2024zhf , and Skyrmions Canfora:2025roy . In what follows, we construct Euclidean AdS wormholes in the presence of non-Abelian ModMax dyons and conformally coupled scalar fields.
@@
$\text{d}s^{2}=\text{d}r^{2}+\frac{h(r)}{4}\left(\sigma_{1}^{2}+\sigma_{2}^{2}+\sigma_{3}^{2}\right)\,.$ (55)
Since the trace of the energy-momentum tensor vanishes on shell by virtue of conformal invariance of scalar and gauge fields, the field equations for the metric yields $R=4\Lambda$. This condition, when evaluated on the ansatz (47), gives the second-order nonlinear equation
@@
This equation can be solved analytically, whose solution is
@@
where $\rho_{0}$ and $r_{0}$ are integration constants with units of length, and the cosmological constant is nonvanishing. In the de Sitter case, the solution is positive definite $\forall r\in\mathbb{R}$ in the range $\rho_{0}\in(-\ell_{\rm dS},\ell_{\rm dS})$, where $\ell_{\rm dS}$ is the de Sitter radius defined as $\Lambda=3/\ell_{\rm dS}^{2}$. In this case, although the solution is regular, the area coordinate exhibits oscillatory behavior rather than a monotonic increase. Therefore, we interpret this solution as a one-parameter extension of the gravitational instanton found in Ref. Hosoya:1989zn .
@@
where $c_{0}$ and $r_{0}$ are integration constants. Notice that, if $c_{0}>0$, this solution is also nowhere vanishing $\forall r\in\mathbb{R}$. Indeed, if $c_{0}=0$, the area function is still positive definite as long as $r_{0}\neq 0$. This solution also represents a wormhole with a vanishing cosmological constant that connects two asymptotically flat regions, with a throat located at $r=0$, where the area coordinate is a minimum.

### v2

@@ 1 Introduction
Recently, it has been realized that it is possible to deform Maxwell theory with a dimensionless coupling $\gamma$, maintaining, at the same time, the electromagnetic duality of the dynamics and the invariance of the action under the conformal group Bandos:2020jsw . The latter implies that if such a continuous deformation exists, it must be controlled by a dimensionless coupling. Such a theory, discovered in Ref. Bandos:2020jsw , was dubbed ModMax. It was later rederived in Ref. Kosyakov:2020wxv using the Gaillard-Zumino-Gibbons-Rasheed formalism for duality rotations in nonlinear electrodynamics Gibbons:1995cv ; Gaillard:1981rj ; Gibbons:1995ap ; Gaillard:1997zr ; Gaillard:1997rt , and more recently in Ref. Ayon-Beato:2024vph employing the Salazar-García-Plebański formulation Salazar:1987ap .
@@
Since ModMax emerges naturally as the most general nonlinear electrodynamics possessing the same features as Maxwell theory, its physical consequences have been thoroughly explored over the last few years. The study of the spectrum of propagating degrees of freedom was done in Ref. Bandos:2020jsw , where it was also shown that there is birrefringence in vacuum. Supersymmetric extensions were constructed in Bandos:2021rqy , while in Flores-Alfonso:2020euz ; Amirabi:2020mzv ; Kubiznak:2022vft ; Barrientos:2022bzm ; Barrientos:2024umq it was shown that its coupling to gravity leads to interesting curved spacetimes as accelerating black holes, multi-black hole solutions Bokulic:2025usc , Euclidean solitons Corral:2025npd , NUTs and Bolts BallonBordo:2020jtw ; Flores-Alfonso:2020nnd , Eguchi-Hanson instantons Colipi-Marchant:2023awk , and hairy black holes Zhang:2021qga ; Ayon-Beato:2024vph . The Galilean and Carrollian limits have been taken and explored in Banerjee:2022sza ; Correa:2024qej (see also Chen:2024vho ). Recently, holographic aspects of ModMax theory have been explored in Ref. Barrientos:2025rde , and their connection to $T\bar{T}$-deformations in conformal field theories was studied in Refs. Rodriguez:2021tcz ; Tempo:2022ndz ; Ferko:2022cix ; Babaei-Aghbolagh:2022leo ; Babaei-Aghbolagh:2022uij ; Ferko:2023wyi ; Babaei-Aghbolagh:2024hti ; Babaei-Aghbolagh:2024uqp .
@@
In Yang-Mills theory, instantons are Euclidean, stationary, and regular (anti-)self-dual solutions that minimize the action on a given topological sector Belavin:1975fg . They also control quantum effects as tunneling between topologically inequivalent vacua, and they are fundamental in a non-perturbative formulation of the quantum theory tHooft:1976rip ; tHooft:1976snw . However, due to topological reasons, Euclidean $U(1)$ gauge theories do not have finite action, localized, regular instantonic solutions on flat space, a feature that is cured by the non-Abelian extensions. Given the relevance of the instantonic configurations, it is natural to ask whether they exist in a non-Abelian extension of the ModMax theory, where the notion of self-duality is natural, and how they are deformed with respect to the instantons in Yang-Mills. This is the aim of the present work.
@@
We start by considering a sensible one-parameter family of non-Abelian extension of ModMax theory in Sec. 2. The theory is parameterized by a single dimensionless constant $\gamma$, which reduces to Yang-Mills theory in the $\gamma\to 0$ limit. We demonstrate that the theory admits (anti-)self-dual configurations with a vanishing energy-momentum tensor, despite the inherent nonlinearities of the ModMax model. Focusing on the $SU(2)$ case, in Sec. 3 we construct the generalized BPST instanton on flat space and extend the solution to Euclidean de Sitter ($\mathbb{S}^{4})$ and anti-de Sitter ($\mathbb{H}^{4})$ backgrounds. In the latter case, the Chern-Pontryagin index is shown to depend on the instanton size, since the configuration is not a pure gauge at infinity; a feature already reported by Callan and Wilczek in Yang-Mills theory on negative-curvature backgrounds Callan:1989em . Section 4 considers the ’t Hooft ansatz, which accommodates multi-instantons in Yang-Mills theory tHooft:1976snw . Indeed, we show that such a choice is also very useful in the non-Abelian ModMax context. Remarkably, the resulting (anti-)self-dual system reduces to a single nonlinear equation, which we solve perturbatively in the dimensionless parameter. To first order, we obtain a formal solution for the $N$-instanton configuration. In Sec. 5, we compute the spectrum of the Dirac operator on the boundary of Euclidean AdS space, which is crucial for obtaining the non-local contributions to the Dirac index in the presence of non-Abelian fields. In Sec. 6, we couple the Euclidean non-Abelian ModMax theory to Einstein gravity in the presence of a conformally-coupled scalar field and construct new gravitating solutions, including Euclidean wormholes and gravitational instantons with a secondary scalar hair. We conclude with comments in Sec. 7.
@@ 2 Non-Abelian ModMax theory
The extension of ModMax electrodynamics to the $SU(2)$ case has been studied in Ref. Cirilo-Lombardo:2023poc . It was shown that the latter can be obtained from the secular equation that arises from the four eigenvectors associated with the four main directions in the Newman-Penrose classification Newman:1961qr . In Euclidean signature, the resulting conformally-invariant action is given by [^1] Here, we will work mainly in Euclidean signature unless otherwise stated. The Lorentzian version of Eq. (4) is obtained by simultaneously changing the global sign of the action and performing $Y^{2}\to-Y^{2}$.
@@
It is worth mentioning that Gibbons and Rasheed studied a global $SO(2)$ electromagnetic duality invariance of general nonlinear electrodynamics in Ref. Gibbons:1995cv . They concluded that the condition $\tilde{F}_{\mu\nu}F^{\mu\nu}=\tilde{P}_{\mu\nu}P^{\mu\nu}$ ensures that the theory is invariant under duality rotations. One can check that the non-Abelian generalization of the Gibbons-Rasheed condition, that is,
@@
is satisfied if the definition in Eq. (6) is assumed. However, Eqs. (5) and (7) are not invariant under global $SO(2)$ duality transformations, as they both depend explicitly on the non-Abelian gauge connection. Hence, due to the nonlinear nature of the field strength, there is no local transformation of the non-Abelian gauge connection that implements a continuous $SO(2)$ rotation.
On the other hand, conformal invariance implies that the stress-energy tensor must be traceless. In the case of the non-Abelian theory, the latter is given by
@@
$F=\frac{1}{2}F_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}=\left(a^{\prime}\text{d}r\wedge\sigma^{i}+\frac{a(a-1)}{2}\epsilon^{i}_{\penalty 10000\ jk}\sigma^{j}\wedge\sigma^{k}\right)t_{i}\,$, (16)
@@
$P=\frac{1}{2}P_{\mu\nu}\text{d}x^{\mu}\wedge\text{d}x^{\nu}$ $=\left(e^{-\gamma}a^{\prime}\text{d}r\wedge\sigma^{i}+\frac{a(a-1)e^{\gamma}}{2}\epsilon^{i}_{\penalty 10000\ jk}\sigma^{j}\wedge\sigma^{k}\right)t_{i}\,$, (18)
@@ 4 Beyond the BPST instanton
where $\Phi=\Phi(x)$ is a function of the coordinates, with no particular symmetry assumed, and $\bar{\eta}_{\mu\nu}:=\bar{\eta}_{a\mu\nu}\,t^{a}$. In the case of Yang-Mills theory, substituting this ansatz into the self-dual condition $F_{\mu\nu}=\tilde{F}_{\mu\nu}$ reduces them to $\nabla^{2}\Phi(x)=0$, which is solved by tHooft:1976snw
@@ 5 Dirac spectrum on the Euclidean AdS boundary
$\lambda_{\pm}$ $=\frac{1}{4}\pm 2\sqrt{\left(p-q-2a_{0\pm}\right)^{2}+4\left(2a_{0\pm}+pq\right)-8a_{0\pm}\sqrt{pq}}\;.$ (43)
Then, the contribution of non-Abelian ModMax instantons to the spectrum of the Dirac operator becomes transparent. This provides an important step towards obtaining the spectral asymmetry of the Dirac operator at the conformal boundary of Euclidean AdS, which is crucial for computing the index of the Dirac operator analytically. If the latter is nonvanishing, it would indicate the presence of axial anomaly induced by self-dual instantons in non-Abelian ModMax theory.
@@ 6.3 Euclidean AdS wormholes and gravitational instantons
Despite their Euclidean nature, they exhibit a wide range of applications in both particle physics and cosmology Hebecker:2018ofv . They also give rise to several conceptual challenges, such as the factorization problem in holography Maldacena:2004rf ; Arkani-Hamed:2007cpn ; Bergshoeff:2004pg ; Andriolo:2022rxc ; Jonas:2023ipa ; Gutperle:2002km . A broad variety of such configurations have been explored in different settings, including supergravity theories Hertog:2017owm ; Anabalon:2018rzq ; Anabalon:2020loe ; Anabalon:2023kcp ; Astesiano:2023iql , higher-curvature extensions of gravity Dotti:2006cp ; Dotti:2007az ; Dotti:2010bw , and in the presence of matter fields such as Yang-Mills fields Maldacena:2004rf ; Hosoya:1989zn , phantom fields Blazquez-Salcedo:2020nsa ; Chew:2021vxh , Higgs fields Betzios:2024zhf , Skyrmions Canfora:2025roy , and with scalar fields Anabalon:2012tu ; Cisterna:2021xxq ; Barrientos:2022avi ; Cisterna:2023uqf . In what follows, we construct Euclidean AdS wormholes in the presence of non-Abelian ModMax dyons and conformally coupled scalar fields.
@@
$\text{d}s^{2}=\text{d}r^{2}+\frac{h(r)}{4}\left(\sigma_{1}^{2}+\sigma_{2}^{2}+\sigma_{3}^{2}\right)\,$, (55)
where $h(r)$ is referred to as the area coordinate. Since the trace of the energy-momentum tensor vanishes on shell by virtue of conformal invariance of scalar and gauge fields, the field equations for the metric yields $R=4\Lambda$. This condition, when evaluated on the ansatz (47), gives the second-order nonlinear equation
@@
This equation can be solved analytically, and its solution is
@@
where $\rho_{0}$ and $r_{0}$ are integration constants of dimension $1$, and the cosmological constant is nonvanishing. In the de Sitter case, the solution is positive definite $\forall r\in\mathbb{R}$ in the range $\rho_{0}\in(-\ell_{\rm dS},\ell_{\rm dS})$, where $\ell_{\rm dS}$ is the de Sitter radius defined as $\Lambda=3/\ell_{\rm dS}^{2}$. In this case, although the solution is regular, the area coordinate exhibits oscillatory behavior rather than a monotonic increase. Therefore, we interpret this solution as a one-parameter extension of the gravitational instanton found in Ref. Hosoya:1989zn .
@@
where $c_{0}$ and $r_{0}$ are integration constants. Notice that, if $c_{0}>0$, this solution is also nowhere vanishing $\forall r\in\mathbb{R}$. Indeed, if $c_{0}=0$, the area coordinate is still positive definite as long as $r_{0}\neq 0$. This solution also represents a wormhole with a vanishing cosmological constant that connects two asymptotically flat regions, with a throat located at $r=0$, where the area coordinate is a minimum.